OpenStudy (anonymous):

Peter is planting trees in his yard. He wants the ratio of pine trees to oak trees to be 6 to 2. Peter wants to plant a total of 72 trees. How many pine trees should he plant?

OpenStudy (phi):

if he plants 6 oak and 2 pine, then you should notice two things: he planted a ratio of \[ \frac{6 \ oak}{2 \ pine} \] which btw simplifes to 3/1 also, he planted a total of 8 trees

OpenStudy (phi):

know the total is 8, you can make up other ratios: \[ \frac{num\ of \ oak}{total} = \frac{6}{8} \] or \[ \frac{num\ of \ pine}{total} = \frac{2}{8} \]

OpenStudy (phi):

now try to fit a total of 72 trees into the last ratio. where does the 72 go ?

OpenStudy (anonymous):

umm idk im not very good at these

OpenStudy (phi):

is 72 the number of pine trees, the number of oak trees or the total number of trees?

OpenStudy (anonymous):

the total number of trees

OpenStudy (phi):

yes, 72 is the total number of trees the question asks How many pine trees should he plant? I think it's obvious we don't know that number. So call it x for "unknown" can you write the fraction number of pine trees divided by total number of trees ? just the fraction?